# Fermionic Averaged Circuit Eigenvalue Sampling

### v1

@@ I Introduction
Noise tailoring is well-suited the setting of gate-dependent noise, where we assume each distinct physical operation on the experiment introduces a unique, reproducible noise channel. The general problem of estimating all noise in a gate set having gate-dependent noise without any averaging is quite challenging [6, 7, 8], so it is natural to ask what role noise tailoring can play to simplify the problem. Ref. [9] introduced averaged circuit eigenvalue sampling (ACES) for simultaneously characterizing the gate-dependent noise of many noisy Clifford gates with Pauli twirling. Clifford circuits with Pauli noise preserve the set of Pauli observables up to an overall factor—called the circuit eigenvalue—which quantifies the noise in the circuit. The main idea behind ACES is that the eigenvalues of many different noisy circuits are all dependent on a common set of gate-level noise eigenvalues, which can be extracted with access to enough experimental data.
@@ A FACES model on $2n$ fermionic modes is specified by a 3-tuple
=($\mathcal{M}=(\{G_{1},\ldots,G_{K}\},\{\boldsymbol{\xi}_{1},\dots,\boldsymbol{\xi}_{K}\},\{C_{1},\ldots,C_{J}\})$ (1)
for $K$ gates $G_{k}$, each of which is associated to a vector of $12n+$ 1 channel eigenvalues $\boldsymbol{\xi}_{k}$, and $J\geq K$ FLO circuits $C_{j}$. From these, we compute the design matrix $\mathbf{A}$, whose integer-valued element $A_{jk}$ specifies the number of occurrences of gate $G_{k}$ in circuit $C_{j}$. A FACES experiment constructs this matrix and samples each circuit $S$ times each. We say that a model is realizable by our experiment if the noisy gates are described by a set of true values for $\boldsymbol{\xi}\coloneqq\cup_{k}\ \boldsymbol{\xi}_{k}$.
@@
Result 2 (Informal Statement of theorem 13).
@@
where $\mathbf{A}^{+}$ is the pseudoinverse of the design matrix, with $S=O\bigl{(}\frac{n\log(m)}{\varepsilon^{2}}\bigr{)}$ samples per circuit.
[…]
The characterization of noisy fermionic linear optical gates is well-studied in the literature [13, 14, 15, 16]. Of particular relevance to our works are the results of refs. [17] and[18]. In ref. [17], the authors propose a method for estimating the average fidelity of noisy matchgate circuits. In [18], the authors introduce a corresponding protocol that does not require an average over the matchgate group. Our result extends the former result to the in situ case under similar physical requirements, though it still requires a group average. Specifically, we require the ability to randomly sample unitaries to perform the FLO twirling. We also require that we are able to reliably prepare the all-zeros state $\left|\mathbf{0}\right\rangle$ and all-plus state $\left|+\right\rangle^{\otimes n}$, and that we generate a set of noisy FLO circuits that act as a particular target unitary on-net (either the identity, or a particular unitary $U_{+}$, to be defined). We furthermore need to be able to measure in the computational basis and the Pauli $Y$-basis on at-least one qubit.
@@
Our paper is organized as follows: in section II, we detail the relevant background on fermionic linear optical circuits (section II.1), Kravchuk matrices (section II.2), twirling (section II.3), and Pauli channel eigenvalues (section II.4). In section III, we explain our measurement protocol, including our assumptions (section III.1 and section III.2), stages of the protocol (section III.3) and sample complexity (section III.4). In particular, we include a proof of our sample complexity guarantee in the main text. Finally, we provide a numerical demonstration of our protocol in section IV and conclude in section V. Our protocol in pseudocode format is detailed in appendix A, and remaining proofs are given in appendix B.
II Background and Notation
II.1 Fermionic Linear Optics
@@
$\{\gamma_{\mu},\gamma_{\nu}\}=2\delta_{\mu\nu}I\mathrm{,}\mbox{\hskip 14.22636pt}\mathrm{and}\mbox{\hskip 14.22636pt}\gamma_{\mu}^{\dagger}=\gamma_{\mu}^{\vphantom{\dagger}}$ (3)
for all $\mu$, $\nu\in[2n]$. Denote the associated group of FLO circuits as $\mathds{F}(2n)$. The elements of this group are Gaussian unitaries up to reflection $U\coloneqq\gamma_{1}^{x}\exp(i\sum_{\mu,\nu}h_{\mu\nu}\gamma_{\mu}\gamma_{\nu})$ for $x\in\{0,1\}$ and $\mathbf{h}$ a real antisymmetric matrix. These preserve linear combinations of Majorana modes under conjugation. Writing $\mathcal{U}(A)\equiv UAU^{\dagger}$ as the unitary conjugation channel by an element $U\in\mathds{F}(2n)$, we have
@@
for $\mathbf{R}=\mathbf{R}^{x}_{1}e^{-4\mathbf{h}}\in O(2n)$ the single-particle transition matrix and $\mathbf{R}_{1}$ the canonical reflection with elements $(\mathbf{R}_{1})_{\mu\nu}=-(-1)^{\delta_{\mu,1}}\delta_{\mu\nu}$. By the canonical anticommutation relations eq. 3, FLO circuits also preserve fixed-degree polynomials of Majorana operators. For a subset $\boldsymbol{\alpha}\subseteq[2n]$, let $\gamma_{\boldsymbol{\alpha}}\equiv\prod_{\mu\in\boldsymbol{\alpha}}\gamma_{\mu}$ denote the Majorana monomial labeled by $\boldsymbol{\alpha}$, where the order of the product is taken such that the index is ascending from left to right. Eq. (4) generalizes as
@@
Linear transformations of the form in eq. 5 constitute a representation of $O(2n)$ by the Cauchy-Binet formula, and this representation is irreducible. As explained in Ref. [17], FLO transformations are uniquely determined up to an overall phase by their corresponding element of $O(2n)$.
It is convenient to generalize our index-subset notation in eq. 6 to binary strings with a fixed Hamming weight. We refer to a binary string by a bold alphabetic character, and, using the same notation as in eq. 6, we denote $\mathbf{x}|k,\ell\rightarrow\{\mathbf{x}\subseteq\mathds{Z}_{2}^{\ell}||\mathbf{x}|=k\}$. In both use cases, we drop explicit reference to $\ell$ where clear from context and simply write “$\boldsymbol{\alpha}|k$” or “$\mathbf{x}|k$”.
@@
$\begin{cases}\gamma_{2j-1}=Z^{\otimes(j-1)}\otimes X_{j}&\\ \gamma_{2j}=Z^{\otimes(j-1)}\otimes Y_{j}&\end{cases}\mathrm{,}$ (7)
where the notation $[n]$ means the set of integers $\{1,\ldots,n\}$. These operators satisfy the canonical anticommutation relations eq. 3. In this context, the group of FLO circuits is isomorphic to the group of matchgate circuits, which we denote as $\mathds{M}(n)$. Matchgate circuits can be decomposed into 2-qubit matchgates of the form
@@ II.2 Kravchuk Matrices
(Kravchuk matrix) The Kravchuk matrix of order $\ell$, denoted by $\mathbf{M}^{(\ell)}\in\mathds{Z}^{(\ell+1)\times(\ell+1)}$, is an integer-valued matrix whose elements are symmetric polynomials evaluated on ±$\pm 1$.
@@
The Walsh-Hadamard matrix of size $2^{n}$ has matrix elements $(-1)^{\mathbf{x}\cdot\mathbf{y}}$ in the standard basis, where $\mathbf{x}$ and $\mathbf{y}$ are bit-strings of length $n$ that label matrix elements. The elements of the Kravchuk matrix are obtained by summing $\sum_{\mathbf{y}|k}(-1)^{\mathbf{x}\cdot\mathbf{y}}$ to get a matrix element that depends only on the Hamming weight $k$ and the bit-string $\mathbf{x}$. In fact, this expression depends only on the Hamming weight of $\mathbf{x}$. By keeping one copy of the resulting sum for each value that $|\mathbf{x}|$ can take, we get the Kravchuk matrix elements in eq. 9.
@@
The proof is the content of section B.1. ∎
@@
[…]
@@
The proof is the content of section B.2. ∎
@@
II.3 Noise Channels and Twirling
@@
where we employ the unitary channel notation from section II.1, and $\mathbb{E}_{x}[f(x)]$ denotes the expectation value over $x$ of $f(x)$. The main examples we consider are Pauli and FLO twirling, where the expectation is given respectively by
@@
[…]
where eq. 21 is a Haar average over the group of FLO unitaries. Note also that twirling is idempotent. That is, twirling twice is the same as twirling once. Additionally, $\mathcal{E}^{\mathcal{S}}$ commutes with every element of $\mathcal{S}$ when $\mathcal{S}$ is a group, by construction.
A channel that is invariant under Pauli twirling is called a Pauli channel. It is easy to show that Pauli twirling a general channel results in a Pauli channel [19, Lemma 5.2.4]. Pauli channels can be parameterized as
@@
where the $\{p_{\mathbf{x}}\}_{\mathbf{x}\in{\sf{P}}(n)}$ form a probability distribution over bit-strings of length $2n$. These are called the Pauli error probabilities. Starting from a general channel with Kraus operators $K_{j}\equiv\sum_{j}c_{j\mathbf{x}}\sigma_{\mathbf{x}}$ as in eq. 17, the relationship to the Pauli error probabilities of the twirled channel in eq. 22 is [19, Lemma 5.2.4]
@@
See section B.3. ∎
We remark that the FLO twirl of a general channel has fermionic error probabilities that are related to the Kraus operators by concatenating the expressions in eq. 23 and eq. 24.
@@
We will eventually require that $\mathcal{E}$, so defined, is not too noisy in a sense made precise in section III.1. Prepending a random FLO unitary $\mathcal{U}^{\dagger}$ to $\widetilde{\mathcal{V}}$ and appending the unitary $\mathcal{U}$ conjugated by $\mathcal{V}$ gives the twirled noise map followed by the ideal channel in expectation. The following statement of this is a standard result.
@@
[…]
@@ II.4 Channel Eigenvalues
$\mathcal{E}(\sigma_{\mathbf{x}})=\lambda_{\mathbf{x}}\sigma_{\mathbf{x}}\,.$ (28)
@@
Now let $\mathcal{E}^{\mathds{F}(2n)}$ be a FLO-twirled channel. Then there is a similar eigenvalue relationship where the eigenvectors are Majorana monomials. From the parameterization in eq. 24, we can derive the eigenvalue relation
@@
=$=\left\{\sum_{k=0}^{2n}\left[\sum_{\boldsymbol{\alpha}|k}(-1)^{|\boldsymbol{\beta}|k+|\boldsymbol{\alpha}\cap\boldsymbol{\beta}|}\right]\binom{2n}{k}^{-1}q_{k}\right\}\gamma_{\boldsymbol{\beta}}\,.$
The coefficient in braces can be simplified using lemma 4 and the definitions in section II.2 to give the FLO-twirled channel eigenvalue
@@
As $2^{-\ell/2}\left[\mathbf{s}^{(\ell)}\cdot\mathbf{M}^{(\ell)}\right]$ is an involution, we can invert this relation and express the probabilities $\mathbf{q}$ from eq. 33 in terms of the eigenvalues $\boldsymbol{\xi}$ as
@@
$\widetilde{\mathcal{U}}^{\mathds{F}(2n)}(\gamma_{\boldsymbol{\beta}})=\left(\prod_{j}\xi_{g,k}\right)\mathcal{U}(\gamma_{\boldsymbol{\beta}})\mbox{\hskip 14.22636pt}(k\coloneqq|\boldsymbol{\beta}|)$ (36)

### v2

@@ 1 Introduction
Noise tailoring is well-suited to the setting of gate-dependent noise, where we assume each distinct physical operation on the experiment introduces a unique, reproducible noise channel. The general problem of estimating all noise in a gate set having gate-dependent noise without any averaging is quite challenging [6, 7, 8], so it is natural to ask what role noise tailoring can play to simplify the problem. Ref. [9] introduced averaged circuit eigenvalue sampling (ACES) for simultaneously characterizing the gate-dependent noise of many noisy Clifford gates with Pauli twirling. Clifford circuits with Pauli noise preserve the set of Pauli observables up to an overall factor—called the circuit eigenvalue—which quantifies the noise in the circuit. The main idea behind ACES is that the eigenvalues of many different noisy circuits are all dependent on a common set of gate-level noise eigenvalues, which can be extracted with access to enough experimental data.
@@ A FACES model on $2n$ fermionic modes is specified by a 3-tuple
$\mathcal{M}=(\{G_{1},\ldots,G_{K}\},\{\boldsymbol{\xi}_{1},\dots,\boldsymbol{\xi}_{K}\},\{C_{1},\ldots,C_{J}\})$ (1)
for $K$ gates $G_{k}$, each of which is associated to a vector of $2n+1$ channel eigenvalues $\boldsymbol{\xi}_{k}$, and $J\geq K$ FLO circuits $C_{j}$. From these, we compute the design matrix $\mathbf{A}$, whose integer-valued element $A_{jk}$ specifies the number of occurrences of gate $G_{k}$ in circuit $C_{j}$. A FACES experiment constructs this matrix and samples each circuit $S$ times each. We say that a model is realizable by our experiment if the noisy gates are described by a set of true values for $\boldsymbol{\xi}\coloneqq\cup_{k}\ \boldsymbol{\xi}_{k}$.
@@
Result 2 (Informal Statement of theorem˜13).
@@
where $\mathbf{A}^{+}$ is the pseudoinverse of the design matrix, with $S=O\bigl(\frac{n\log(J)}{\varepsilon^{2}}\bigr)$ samples per circuit.
The norm on $\mathbf{A}^{+}$ is the induced $\infty\to\infty$ norm, defined as $\lVert X\rVert_{\infty}=\mathrm{max}_{i}\sum_{j}|X_{ij}|$. In the setting where $\mathbf{A}$ does not have full column-rank, its psuedoinverse is still well-defined, but the model includes at least one unlearnable, or “gauge” parameter that is inaccessible to our framework (for a detailed investigation of this, see [11, 12]).
Considering eq. (2), one imagines minimizing $\rVert\mathbf{A}^{+}\lVert_{\infty}$ as a promising method to likely improve the precision of the estimate. Previous work has considered optimizing gate-dependent noise learning in a restricted setting (this is done in gate-set tomography by amplifying certain errors using long sequences [8], for example). Other methods have also been explored in an ACES-type context. For example, the authors of ref. [13] consider optimizing a figure of merit related to the gate-eigenvalue estimator covariance matrix. A comprehensive theory is lacking and currently constitutes an open research direction.
The characterization of noisy fermionic linear optical gates is well-studied in the literature [14, 15, 16, 17]. Of particular relevance to our works are the results of refs. [18] and [19]. In ref. [18], the authors propose a method for estimating the average fidelity of noisy matchgate circuits. In [19], the authors introduce a corresponding protocol that does not require an average over the matchgate group. Our result extends the former result to the in situ case under similar physical requirements, though it still requires a group average. Specifically, we require the ability to randomly sample unitaries to perform the FLO twirling and that the effect of the noise on the twirling gates can itself be incorporated into our framework. We return to this point throughout the paper.
We also require that we are able to reliably prepare the all-zeros state $\left|\mathbf{0}\right\rangle$ and all-plus state $\left|+\right\rangle^{\otimes n}$, and that we generate a set of noisy FLO circuits that act as a particular target unitary on-net (either the identity, or a particular unitary $U_{+}$, to be defined). We furthermore need to be able to measure in the computational basis and the Pauli $Y$-basis on at least one qubit.
@@
Our paper is organized as follows: in section˜2, we detail the relevant background on fermionic linear optical circuits (section˜2.1), Kravchuk matrices (section˜2.2), twirling (section˜2.3), and Pauli channel eigenvalues (section˜2.4). In section˜3, we explain our measurement protocol, including our assumptions (section˜3.1 and section˜3.2), stages of the protocol (section˜3.3) and sample complexity (section˜3.4). In particular, we include a proof of our sample complexity guarantee in the main text. Finally, we provide a numerical demonstration of our protocol in section˜4 and conclude in section˜5. Our protocol in pseudocode format is detailed in appendix˜A, and remaining proofs are given in appendix˜B.
2 Background and Notation
2.1 Fermionic Linear Optics
@@
[…]
for all $\mu$, $\nu\in[2n]$, where the notation $[n]$ means the set of integers $\{1,\ldots,n\}$. Denote the associated group of FLO circuits as $\mathds{F}(2n)$. The elements of this group are Gaussian unitaries up to reflection $U\coloneqq\gamma_{1}^{x}\exp(\sum_{\mu,\nu}h_{\mu\nu}\gamma_{\mu}\gamma_{\nu})$ for $x\in\{0,1\}$ and $\mathbf{h}$ a real antisymmetric matrix. These preserve linear combinations of Majorana modes under conjugation. Writing $\mathcal{U}(A)\equiv UAU^{\dagger}$ as the unitary conjugation channel by an element $U\in\mathds{F}(2n)$, we have
@@
for $\mathbf{R}=\mathbf{R}^{x}_{1}e^{-4\mathbf{h}}\in O(2n)$ the single-particle transition matrix and $\mathbf{R}_{1}$ the canonical reflection with elements $(\mathbf{R}_{1})_{\mu\nu}=-(-1)^{\delta_{\mu,1}}\delta_{\mu\nu}$. By the canonical anticommutation relations, eq.˜3, FLO circuits also preserve fixed-degree polynomials of Majorana operators. For a subset $\boldsymbol{\alpha}\subseteq[2n]$, let $\gamma_{\boldsymbol{\alpha}}\equiv\prod_{\mu\in\boldsymbol{\alpha}}\gamma_{\mu}$ denote the Majorana monomial labeled by $\boldsymbol{\alpha}$, where the order of the product is taken such that the index is ascending from left to right. Eq. (4) generalizes as
@@
Linear transformations of the form in eq.˜5 constitute a representation of $O(2n)$ by the Cauchy-Binet formula, and this representation is irreducible. As explained in ref. [18], FLO transformations are uniquely determined up to an overall phase by their corresponding element of $O(2n)$.
It is convenient to generalize our index-subset notation in eq.˜6 to binary strings with a fixed Hamming weight. We refer to a binary string by a bold alphabetic character, and, using the same notation as in eq.˜6, we denote $\mathbf{x}|k,\ell\rightarrow\{\mathbf{x}\subseteq\mathds{Z}_{2}^{\ell}||\mathbf{x}|=k\}$. In both use cases, we drop explicit reference to $\ell$ where clear from context and simply write $\boldsymbol{\alpha}|k$” or $\mathbf{x}|k$”.
@@
$\begin{cases}\gamma_{2j-1}=Z^{\otimes(j-1)}\otimes X_{j}&\\ \gamma_{2j}=Z^{\otimes(j-1)}\otimes Y_{j}&\end{cases}\mathrm{.}$ (7)
These operators satisfy the canonical anticommutation relations eq.˜3. In this context, the group of FLO circuits is isomorphic to the group of matchgate circuits, which we denote as $\mathds{M}(n)$. Matchgate circuits can be decomposed into 2-qubit matchgates of the form
@@ 2.2 Kravchuk Matrices
(Kravchuk matrix) The Kravchuk matrix of order $\ell$, denoted by $\mathbf{M}^{(\ell)}\in\mathds{Z}^{(\ell+1)\times(\ell+1)}$, is an integer-valued matrix whose elements are symmetric polynomials evaluated on $\pm 1$.
@@
The Walsh-Hadamard matrix of size $2^{n}$ has matrix elements $(-1)^{\mathbf{x}\cdot\mathbf{y}}$ in the standard basis, where $\mathbf{x}$ and $\mathbf{y}$ are bit-strings of length $n$ that label matrix elements. The elements of the Kravchuk matrix are obtained by summing $\sum_{\mathbf{y}|k}(-1)^{\mathbf{x}\cdot\mathbf{y}}$ to get a matrix element that depends only on the Hamming weight $k$ and the bit-string $\mathbf{x}$. In fact, this expression depends only on the Hamming weight of $\mathbf{x}$. By keeping one copy of the resulting sum for each value that $|\mathbf{x}|$ can take, we get the Kravchuk matrix elements in eq.˜9.
@@
The proof is the content of section˜B.1. ∎
@@
$s^{(\ell)}_{jk}=\delta_{jk}\delta_{j\ (\mathrm{mod}\ 2),0}+\delta_{\ell-j,k}\delta_{j\ (\mathrm{mod}\ 2),1}\mathrm{.}$ (14)
@@
The proof is the content of section˜B.2. ∎
@@
2.3 Noise Channels and Twirling
@@
where we employ the unitary channel notation from section˜2.1, and $\mathbb{E}_{x}[f(x)]$ denotes the expectation value over $x$ of $f(x)$. The main examples we consider are Pauli and FLO twirling, where the expectation is given respectively by
@@
$\mathcal{E}^{\mathds{F}(2n)}(X)$ $=\int_{U\in\mathds{F}(2n)}dU\ \sum_{j}\mathcal{U}\left(K_{j}\right)X\ \mathcal{U}\left(K_{j}\right)^{\dagger}$ (21)
where eq.˜21 is a Haar average over the group of FLO unitaries. Note also that twirling is idempotent. That is, twirling twice is the same as twirling once. Additionally, $\mathcal{E}^{\mathcal{S}}$ commutes with every element of $\mathcal{S}$ when $\mathcal{S}$ is a group, by construction.
A channel that is invariant under Pauli twirling is called a Pauli channel. It is easy to show that Pauli twirling a general channel results in a Pauli channel [20, Lemma 5.2.4]. Pauli channels can be parameterized as
@@
where the $\{p_{\mathbf{x}}\}_{\mathbf{x}\in{\sf{P}}(n)}$ form a probability distribution over bit-strings of length $2n$. These are called the Pauli error probabilities. Starting from a general channel with Kraus operators $K_{j}\equiv\sum_{j}c_{j\mathbf{x}}\sigma_{\mathbf{x}}$ as in eq.˜17, the relationship to the Pauli error probabilities of the twirled channel in eq.˜22 is [20, Lemma 5.2.4]
@@
See section˜B.3. ∎
We remark that the FLO twirl of a general channel has fermionic error probabilities that are related to the Kraus operators by concatenating the expressions in eqs.˜23, 2 and 25.
@@
We will eventually require that $\mathcal{E}$, so defined, is not too noisy in a sense made precise in section˜3.1. Prepending a random FLO unitary $\mathcal{U}^{\dagger}$ to $\widetilde{\mathcal{V}}$ and appending the unitary $\mathcal{U}$ conjugated by $\mathcal{V}$ gives the twirled noise map followed by the ideal channel in expectation. The following statement of this is a standard result.
@@
$\widetilde{\mathcal{V}}^{\mathcal{S}}\coloneqq\int_{\mathcal{S}}d\mathcal{U}\left(\mathcal{V}\mathcal{U}\mathcal{V}^{\dagger}\right)\left(\mathcal{V}\mathcal{E}\right)\mathcal{U}^{\dagger}=\mathcal{V}\mathcal{E}^{\mathcal{S}}\,.$ (27)
@@ 2.4 Channel Eigenvalues
$\mathcal{E}^{\mathds{P}(n)}(\sigma_{\mathbf{x}})=\lambda_{\mathbf{x}}\sigma_{\mathbf{x}}\,.$ (28)
@@
Now let $\mathcal{E}^{\mathds{F}(2n)}$ be a FLO-twirled channel. Then there is a similar eigenvalue relationship where the eigenvectors are Majorana monomials. From the parameterization in eq.˜24, we can derive the eigenvalue relation
@@
$=\left\{\sum_{k=0}^{2n}\left[\sum_{\boldsymbol{\alpha}|k}(-1)^{|\boldsymbol{\beta}|k+|\boldsymbol{\alpha}\cap\boldsymbol{\beta}|}\right]\binom{2n}{k}^{-1}q_{k}\right\}\gamma_{\boldsymbol{\beta}}\,.$
The coefficient in braces can be simplified using lemma˜4 and the definitions in section˜2.2 to give the FLO-twirled channel eigenvalue
@@
As $2^{-\ell/2}\left[\mathbf{s}^{(\ell)}\cdot\mathbf{M}^{(\ell)}\right]$ is an involution, we can invert this relation and express the probabilities $\mathbf{q}$ from eq.˜33 in terms of the eigenvalues $\boldsymbol{\xi}$ as
@@
$\widetilde{\mathcal{U}}^{\mathds{F}(2n)}(\gamma_{\boldsymbol{\beta}})=\left(\prod_{g}\xi_{g,k}\right)\mathcal{U}(\gamma_{\boldsymbol{\beta}})\mbox{\hskip 14.22636pt}(k\coloneqq|\boldsymbol{\beta}|)$ (36)